Answer? write the equation of the line that is parallel to the given line and passes through the given point. y= -2x-1, (2,6)
Answer by mananth(16949)

y =-2 x-1
Compare this equation with y=mx+b
slope m = -2
The slope of a line parallel to the above line will be the same
The slope of the required line will be -2
m= -2 ,point (2,6 )
Find b by plugging the values of m & the point in
y=mx+b
6 = -4 + b
b= 10
m= -2
Plug value of the slope and b in y = mx +b
The required equation is
y = -2x+ 10
m.ananth@hotmail.ca